L.J.’s Toys Inc. just purchased a $510,000 machine to produce toy cars. The machine will be fully depreciated by the straight-line method over its six-year economic life. Each toy sells for $27. The variable cost per toy is $12, and the firm incurs fixed costs of $287,000 each year. The corporate tax rate for the company is 35 percent. The appropriate discount rate is 11 percent. What is the financial break-even point for the project? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

L.J.’s Toys Inc. just purchased a $510,000 machine to produce toy cars. The machine will be fully depreciated by the straight-line method over its six-year economic life. Each toy sells for $27. The variable cost per toy is $12, and the firm incurs fixed costs of $287,000 each year. The corporate tax rate for the company is 35 percent. The appropriate discount rate is 11 percent. What is the financial break-even point for the project? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

At financial breakeven point, NPV is zero.

NPV= PV of cash flows – initial investment

0 = OCF x PVIFA (n,R) – 510,000

OCF x PVIFA (n,R) = 510,000

OCF x PVIFA (6,11%)=510,000

OCF = 510,000/ 4.230537

OCF= 120,552.07

Now we need to compute quantity to arrive at this ocf.

Annual depreciation = 510,000/6

                                         = 85,000

Depreciation tax benefit = depreciation x tax rate

                                                = 85000 x 35%

                                                = 29,750

OCF = [(price – VC) x Q - FC] x(1- tax rate) + Depreciation tax benefit

120,552.07 = [(27-12) x Q -287,000](1-0.35) +29750

9.75Q = 120,552.07 -29750+ 186,550

Q= 277,352.07/9.75

Q=28,446.37

Hence, financial break even level of quantity would be 28,446 units.

Thanks

Let me know if you have any doubt